### v1

# Multi-Channel Hypergraph Contrastive Learning for Matrix Completion

Abstract.
Rating is a typical user explicit feedback that visually reflects how much a user likes a related item. The (rating) matrix completion is essentially a rating prediction process, which is also a significant problem in recommender systems. Recently, graph neural networks (GNNs) have been widely used in matrix completion, which captures users’ preferences over items by formulating a rating matrix as a bipartite graph. However, existing methods are susceptible due to data sparsity and long-tail distribution in real-world scenarios. Moreover, the messaging mechanism of GNNs makes it difficult to capture high-order correlations and constraints between nodes, which are essentially useful in recommendation tasks. To tackle these challenges, we propose a Multi-Channel Hypergraph Contrastive Learning framework for matrix completion, named MHCL. Specifically, MHCL adaptively learns hypergraph structures to capture high-order correlations between nodes and jointly captures local and global collaborative relationships through attention-based cross-view aggregation. Additionally, to consider the magnitude and order information of ratings, we treat different rating subgraphs as different channels, encourage alignment between adjacent ratings, and further achieve the mutual enhancement between different ratings through multi-channel cross-rating contrastive learning. Extensive experiments on five public datasets demonstrate that the proposed method significantly outperforms the current state-of-the-art approaches.
Graph Representation Learning, Matrix Completion, Hypergraph Contrastive Learning, Recommendation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Recently, the exponential growth of information has made matrix completion essential for e-commerce and social media platforms (Zhang et al., 2019; Ricci et al., 2010; Xia et al., 2020). These systems typically process two types of user feedback: explicit and implicit (Chen and Wang, 2022). Explicit feedback directly captures user preferences—for example, ratings on a matrix, where each entry quantifies how much a user likes a specific item. In contrast, implicit feedback indirectly reflects preferences, typically inferred as positive signals. This includes behavioral data such as click records, items added to shopping carts, and user favorites (Palomares et al., 2018; Raza and Ding, 2019). Matrix completion is a fundamental problem with various applications, ranging from missing traffic data imputation, protein relation prediction, image restoration, to recommender systems. Specifically, in recommender systems, the aim is to predict missing entries in a partially observed user-item matrix using available explicit or implicit feedback.
Matrix factorization (Candes and Recht, 2012; Kalofolias et al., 2014; Xu et al., 2013) is a traditional approach to matrix completion that decomposes the user-item rating matrix into latent factors representing users and items. In recent years, GNNs have gained prominence in recommender systems (Monti et al., 2017; Shen et al., 2021; Zhang and Chen, 2019; Zhang et al., 2022). These methods model the rating matrix as a bipartite graph, reformulating the matrix completion problem as a link prediction task on this graph. Inductive matrix completion methods, such as IGMC (Zhang and Chen, 2019), utilize GNNs to learn local graph patterns from 1-hop subgraphs around target edges, enabling generalization to unseen nodes. Building on this idea, GIMC (Zhang et al., 2022) introduces edge embeddings and hyperbolic geometry to further enhance the modeling of bipartite graphs. Moreover, some review-based models (Wu et al., 2019; Gao et al., 2020a; Shuai et al., 2022) leverage auxiliary review information to enrich the learning of embeddings, complementing the structure-based methods.
Despite their remarkable success, existing GNN-based matrix completion methods face several key challenges. First, existing GNN-based methods struggle to capture high-order correlations and complex constraints due to their reliance on existing edges in bipartite graphs, which often require stacking multiple layers and can introduce noise, leading to over-smoothing issues (Min et al., 2020). Second, these methods are inherently vulnerable to data sparsity and long-tail distributions (Singh, 2020), where a few highly active nodes dominate interactions, resulting in popularity bias and diminished effectiveness for less active nodes. Many approaches also overlook the magnitude and ordinal nature of ratings. For example, a rating of ‘4’ is closer to ‘5’ than to ‘1’. Although some recent methods (Zhang and Chen, 2019; Shen et al., 2021) incorporate regularization to address certain limitations, they often fail to fully account for these aspects within the model design.
Motivated by the need to address the aforementioned challenges, we propose a novel Multi-Channel Hypergraph Contrastive Learning (MHCL) framework for matrix completion. First, we partition the bipartite graph into multiple subgraphs based on different types of rating edges using a subgraph partition and graph convolution module. During training, an adaptive hypergraph structure learning module dynamically learns hypergraph structures for these subgraphs at different layers, capturing high-order correlations between nodes to overcome the first challenge. Second, we introduce a multi-channel cross-rating contrastive learning module that treats each rating subgraph as a distinct channel, promoting mutual enhancement by aligning adjacent views. Third, to address the data sparsity and long-tail distribution challenge, we design an attention-based cross-view fusion module that strengthens interactions between different rating channels. Finally, an attention-based cross-view aggregation module learns node representations collaboratively from both local and global perspectives, fusing them into the final representation. MHCL demonstrates significant performance improvements, achieving up to +9.04% lift in MSE over SOTA baselines for matrix completion and a +10.31% increase in NDCG@10 for recommendation tasks on the ML-100K dataset.
To sum up, this paper makes the contributions as follows:
  • MHCL leverages hypergraph learning to partition the user-item relationship graph into distinct subgraphs, representing various types of rating edges. It dynamically learns hypergraph structures to capture higher-order relationships during training.
  • MHCL introduces a multi-channel cross-rating contrastive learning, treating different rating subgraphs as unique perspectives. This approach promotes alignment between adjacent perspectives, enhancing the modeling of interrelationships between diverse rating types and improving prediction accuracy.
  • MHCL includes an attention-based cross-view aggregation module to integrate information from various rating subgraphs. It employs contrastive learning to collaboratively learn node representations from both local and global perspectives, ultimately fusing them to obtain the final node representations. This multi-view approach comprehensively captures node features, making contributions to improving recommendations.
  • We conduct comprehensive experiments to evaluate the performance of our MHCL model on matrix completion and recommendation tasks, along with diverse baselines, using multiple benchmark datasets.
2. Related Work
2.1. GNNs for Matrix Completion
Recent advancements in graph neural networks (GNNs) have significantly enhanced matrix completion tasks. GNN-based methods interpret the matrix as a bipartite graph, where observed ratings or purchases are represented by edges, transforming matrix completion into a graph-based predictive task. Traditional GNN approaches (Kipf and Welling, 2016; Hamilton et al., 2017; Veličković et al., 2017) use multi-hop convolutions to capture high-order relations from user-item graphs. ST-GCN (Yan et al., 2018), NGCF (Wang et al., 2019) and PinSage (Ying et al., 2018) explore high-order connectivity through multiple graph convolutions, while LightGCN (He et al., 2020) emphasizes neighborhood aggregation by omitting non-linear transformations. Inductive methods like IDCF-NN (Wu et al., 2021b) and IDCF-GC (Wu et al., 2021b) utilize matrix factorization and neural information transfer for embedding computation. GC-MC (Berg et al., 2017) refines node representations using a GNN encoder and decoder, and IMC-GAE (Shen et al., 2021) improves generalization with role-aware attributes and one-hot encoding. IGMC (Zhang and Chen, 2019) extracts enclosing subgraphs to learn local patterns, while GIMC (Zhang et al., 2022) addresses data sparsity with hyperbolic regression. RMG (Wu et al., 2019) integrates graph signals with review information, and RGCL (Shuai et al., 2022) enhances this with semantic review features and contrastive learning (CL). MAGCL (Wang et al., 2023) further refines interaction modeling using multi-view representation, graph contrastive learning, and multi-task learning to capture fine-grained semantic information.
In summary, these advancements within the realm of graph neural networks have ushered in transformative approaches to matrix completion, showcasing their potential to revolutionize recommender systems.
2.2. Hypergraph Learning for Recommendation
Inspired by the flexibility of hypergraphs to overcome the constraints of fixed graph structures in message passing (Gao et al., 2020b; Zhou et al., 2006; Bu et al., 2010; Yang et al., 2019), recent advancements in recommender systems have integrated hypergraph connections to improve relational learning. DualHGCN (Xue et al., 2021) converts multiplex bipartite networks into homogeneous hypergraphs and uses spectral hypergraph convolutional operators for effective intra- and inter-domain information exchange. Similarly, MHCN (Yu et al., 2021) utilizes social-based hypergraph convolutional networks to exploit high-order user relations in a multi-channel framework, while DHCF (Ji et al., 2020) focuses on modeling high-order correlations between users and items. However, these methods rely on manually crafted, static heuristic hypergraph structures.
To address this, recent approaches have explored dynamic hypergraph learning. For example, DHSL (Zhang et al., 2018) uses labels as spatial feature vectors for clustering, and DHGNN (Jiang et al., 2019) applies K-NN and K-means clustering to update hypergraph structures based on local and global features. Additionally, HCCF (Xia et al., 2022) and DHLCF (Wei et al., 2022) use collaborative signals and topological information to dynamically determine hyperedge assignments with MLPs. In contrast, our approach dynamically learns a hypergraph for each rating subgraph, capturing global collaborative relationships at a finer level among homogeneous nodes. Moreover, in addition to the hypergraph structure, DGCLR (Ren et al., 2022), VGCL (Yang et al., 2023) and ReHCL (Wang et al., 2024) consider fusing user comments and ratings for recommendation and introduce the user’s emotional tendency into the node representation learning to improve the recommendation performance.
3. Problem Definition
We set $\mathbf{U}=\{u_{1},u_{2},\ldots,u_{M}\}$ as the set of users and $\mathbf{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ as the set of items, where $M$ and $N$ are the number of users and items, respectively. We use a binary matrix $\mathbf{R}$ to store historical ratings, where $r_{uv}$ indicates a rating from user $u$ to item $v$, and $r_{uv}=0$ means item $v$ is unexposed to user $u$. Interaction data is represented as a user-item bipartite graph $G$, and the task is to predict missing ratings in this graph.
Recent developments in graph neural networks have predominantly applied matrix completion techniques to bipartite user-item graphs, treating it as a link prediction problem. This approach utilizes GNNs to capture complex relationships and dependencies within user-item interactions, advancing recommender systems and collaborative filtering approaches.
Key notations are summarized in Table 5 in Supplement.
4. Methodology
Figure 1. The outline of MHCL’s overall architecture.
In this section, we provide a comprehensive overview of the proposed MHCL framework, as illustrated in Figure 1. MHCL comprises four key components: (1) Subgraph Partition and Graph Convolution: MHCL begins by partitioning the bipartite user-item graph into subgraphs based on different ratings. Graph convolutions are applied to these subgraphs to capture collaborative signals and extract local topological features for users and items across various rating contexts. (2) Multi-Channel Cross-Rating Contrastive Learning: To enhance interaction between adjacent rating subgraphs, MHCL employs a multi-channel cross-rating contrastive learning module. This module promotes information exchange between rating channels, addressing data sparsity and mitigating the long-tail distribution problem. (3) Adaptive Hypergraph Structure Learning: Using local embeddings from the subgraphs, MHCL integrates an adaptive hypergraph structure learning module. This module constructs hypergraphs by identifying high-order dependencies among nodes of the same type (e.g., users with users, items with items), capturing global topological features unique to each rating. This process helps mitigate over-smoothing and enhances the model’s ability to learn higher-order correlations. (4) Attention-Based Cross-View Aggregation: MHCL aggregates local and global representations through an attention-based cross-view aggregation module. By maximizing the consistency between embeddings from diverse views, it leverages global-local contrastive learning techniques to obtain robust final representations.
Through these four components, MHCL delivers a comprehensive solution for matrix completion, leveraging collaborative signals, topological features, and higher-order dependencies to enhance the accuracy and efficiency of recommendations.
4.1. Subgraph Partition and Graph Convolution
Inspired by previous models (Shen et al., 2021; Zhang et al., 2022; Zhang and Chen, 2019), we adopt the initialization method from IMC-GAE (Shen et al., 2021). Let $M$ and $N$ represent the number of users and items, respectively, and $\mathcal{|R|}$ be the number of rating categories. The initial node embeddings are parameterized by a learnable matrix $\mathbf{E}\in\mathbb{R}^{\mathcal{|R|}\times(M+N+3)\times d}$. Each node’s initial features consist of three indices: identical index $\mathbf{id}$, role-aware index $\mathbf{ra}$, and one-hot index $\mathbf{oh}$. We set the identical index to 0, the user’s role-aware index to 1, the item’s to 2, and the one-hot index ranges from 3 to ($M$+$N$+3). Using these initial features, we derive the node’s identical, role-aware, and one-hot embeddings from $\mathbf{E}$. Finally, we concatenate the three embeddings as follows:
(1) $x_{r}^{0}[i]=Concat(\mathbf{E}_{r}[\mathbf{id}[i]],\mathbf{E}_{r}[\mathbf{ra}[i]],\mathbf{E}_{r}[\mathbf{oh}[i]]),$
where $\mathbf{E}_{r}\in\mathbb{R}^{(M+N+3)\times d}$ denotes $r$-th trainable parameter matrix in $\mathbf{E}\in\mathbb{R}^{\mathcal{|R|}\times(M+N+3)\times d}$, and $x_{r}^{0}[i]$ denotes the initializing embedding of the $i$-th node in the $r$-th rating subgraph. The node embeddings are the input of the message-passing layer.
Then we systematically partition the graph into distinct rating subgraphs based on the different types of rating edges. For example, with five rating types, we create five rating subgraphs, each containing node embeddings specific to its rating category. MHCL then operates independently on these subgraphs, applying graph convolutions to derive node embeddings within each subgraph. Unlike conventional GCNs, MHCL forgoes the use of a learnable matrix for feature transformation, opting for an explicit propagation process, which can be expressed as follows:
(2) $x^{r}_{l+1}[i]=\sum\limits_{j\in\mathcal{N}^{r}(i)}\frac{1}{\sqrt{|\mathcal{N}^{r}(i)|\cdot|\mathcal{N}^{r}(j)|}}x^{r}_{l}[j],$
where $\mathcal{N}^{r}(i)$ represents the neighboring node set in the $r$-th subgraph, $x^{r}_{l+1}[i]$ denotes the embedding of the $i$-th node at layer $l+1$ in the $r$-th rating subgraph. By performing graph convolution on diverse rating subgraphs, we obtain multiple embeddings of nodes that are specific to different rating types, and these embeddings can represent the collaborative signal and the topological information of each user and item on different subgraphs.
4.2. Multi-Channel Cross-Rating Contrastive Learning
Unlike traditional heterogeneous graphs with multiple edge types (e.g., click, purchase), our scenario focuses on comparable ratings, where adjacent ratings indicate similar user preferences and node representations. However, the long-tail distribution in real-world datasets complicates MHCL’s partitioning into distinct subgraphs based on rating types, risking the loss of magnitude and order information during graph convolution.
To address this, we introduce a multi-channel cross-rating contrastive learning module. Each rating subgraph acts as a separate channel, capturing unique perspectives, while cross-rating contrastive learning identifies cooperative associations. By using the InfoNCE loss (Gutmann and Hyvärinen, 2010), our approach improves model performance by treating embeddings from adjacent rating subgraphs as positive pairs, preserving the rating order:
(3) ${\mathcal{L}_{S(k)}^{\mathbf{U}}=\sum_{u\in\mathbf{U}}-\log\frac{exp((x_{u}^{(k)}\cdot x_{u}^{(k+1)}/\tau))}{\sum_{u^{\prime}\in\mathbf{U}}exp((x_{u^{\prime}}^{(k)}\cdot x_{u^{\prime}}^{(k+1)}/\tau))},}$
where $x_{u}^{(k)}$ is the normalized output embedding of $k$-th rating subgraph and $\tau$ is the temperature hyper-parameter of softmax.
Most real-world datasets typically use 5-point, 10-point, and 100-point Likert scales. For instance, in a 100-point scale, directly performing contrastive learning across all ratings is impractical. To tackle this, we map the task to a 10-point Likert scale, conducting contrastive learning only between embeddings from rating subgraphs that are evenly divisible by 10.
Using a 5-point Likert scale as an example, $L_{S(1)}^{\mathbf{U}}$ denotes the cross-rating contrastive loss between the subgraphs with ratings ‘1’ and ‘2’, while $L_{S(2)}^{\mathbf{U}}$ represents the loss between the subgraphs with ratings ‘2’ and ‘3’. The overall cross-rating contrastive loss for the entire user can be expressed as follows:
(4) $\mathcal{L}_{S}^{\mathbf{U}}=\mathcal{L}_{S(1)}^{\mathbf{U}}+\mathcal{L}_{S(2)}^{\mathbf{U}}+\dots+\mathcal{L}_{S(\mathcal{|R|}-1)}^{\mathbf{U}}.$
where $\mathcal{|R|}$ denotes the number of rating categories. In a similar way, the cross-ratings contrastive loss of the item side $\mathcal{L}_{S}^{\mathbf{V}}$ can be obtained as:
(5) $\mathcal{L}_{S}^{\mathbf{V}}=\mathcal{L}_{S(1)}^{\mathbf{V}}+\mathcal{L}_{S(2)}^{\mathbf{V}}+\dots+\mathcal{L}_{S(\mathcal{|R|}-1)}^{\mathbf{V}},$
and the complete cross-rating contrastive objective function is the sum of the above two losses:
(6) $\mathcal{L}_{S}=\mathcal{L}_{S}^{\mathbf{U}}+\mathcal{L}_{S}^{\mathbf{V}}.$
4.3. Adaptive Hypergraph Structure Learning
The embedding from Eq. (2) captures collaborative signals and topological information within rating subgraphs, reflecting implicit user preferences or item features. To address the limitations of bipartite graphs, where edges connect nodes of the same type, MHCL employs hypergraph structure learning for users and items. By utilizing hypergraph convolution, MHCL captures global information beyond node types. Inspired by previous work (Wei et al., 2022; Xia et al., 2022), we use an MLP to compute hyperedge assignments within each rating subgraph, resulting in hypergraphs $\mathbf{H}^{r}_{u}$ for users and $\mathbf{H}^{r}_{v}$ for items, which capture complex high-order relationships and dependencies.
(7) $\mathbf{H}^{r}_{u}=Norm(LeakyReLU({x_{u}^{r}}\mathbf{W}_{u}^{r})),$
$\quad\mathbf{H}^{r}_{v}=Norm(LeakyReLU({x_{v}^{r}}\mathbf{W}_{v}^{r})),$
where $Norm(\cdot)$ denotes the normalization function, $\mathbf{W}_{u}^{r}\in\mathbb{R}^{d\times K}$ are trainable weight matrices, in which $K$ is the number of hyperedges.
After deriving the user and item hypergraphs for different ratings, we incorporate implicit high-order correlations into the updated representations of users and items. Drawing from traditional hypergraph convolution methods (Feng et al., 2019; Ji et al., 2020), we follow the spectral hypergraph convolution framework. The process begins by aggregating node embeddings to form hyperedge representations via hyperedge convolution. These hyperedge embeddings are then re-aggregated to generate updated node representations. The hypergraph convolution, applied independently to each rating subgraph, is formally defined as follows:
(8) $h_{u}^{r}={\mathbf{H}}_{u}^{r}\mathbf{D}_{u}^{-1}{\mathbf{H}}_{u}^{r\mathsf{T}}{x_{u}^{r}},\quad h_{v}^{r}={\mathbf{H}}_{v}^{r}\mathbf{D}_{v}^{-1}{\mathbf{H}}_{v}^{r\mathsf{T}}{x_{v}^{r}}.$
Here, $\mathbf{D}_{u}$ and $\mathbf{D}_{v}$ represent the vertex degree matrices associated with the learned user hypergraph $\mathbf{H}_{u}^{r}$ and item hypergraph $\mathbf{H}_{v}^{r}$, respectively. These matrices play a crucial role in the re-scaling of the resulting embeddings.
4.4. Attention-Based Cross-View Aggregation
As shown in Eq. (2) and Eq. (8), each node gains a local embedding for the local topology and a global embedding for higher-order dependencies within each rating subgraph. However, downstream tasks require compact embeddings, and traditional pooling methods (e.g., average or sum) often overlook the importance of relations between different node representations.
To overcome this limitation, we propose an attention-based cross-view contrastive aggregation mechanism to reinforce relationships between adjacent ratings. Formally, let $\mathcal{N}_{r}$ denote the two ratings adjacent to rating $r$, and let $x^{r}$ represent the local embeddings of nodes in the $r$-th rating subgraph. The message passing between adjacent ratings is carried out as follows:
(9) $z^{r}=x^{r}+\sum\limits_{r^{\prime}\in\mathcal{N}_{r}}\lambda^{r^{\prime}}\cdot x^{r^{\prime}},$
where $\lambda^{r^{\prime}}$ is the normalized relevance of rating $r^{\prime}$ to rating $r$, and it is calculated by:
(10) $\begin{gathered}\lambda^{r^{\prime}}=\frac{\exp\left(LeakyReLU\left(q^{r}x^{r^{\prime}}\right)\right)}{\sum_{r^{\prime}\in\mathcal{N}_{r}}\exp\left(LeakyReLU\left(q^{r}x^{r^{\prime}}\right)\right)}.\end{gathered}$
The vector $q^{r}$ is a trainable attention vector tailored to the rating $r$. It serves as a pivotal element in regulating the information flow between relation $r$ and the neighboring relations found within $\mathcal{N}_{r}$. Similarly, we employ the attention-based cross-view aggregation mechanism on the global embeddings in a parallel manner:
(11) $\Gamma^{r}=h^{r}+\sum\limits_{r^{\prime}\in\mathcal{N}_{r}}\delta^{r^{\prime}}\cdot h^{r^{\prime}},$
where $h^{r}$ signifies the global embeddings of nodes within the $r$-th rating subgraph. Similar to $\lambda^{r^{\prime}}$ in Eq. (10), $\delta^{r^{\prime}}$ can be calculated according to the similar computation process like Eq. (10). Subsequently, we consolidate all nodes’ global and local embeddings across rating subgraphs into a unified vector representation using a summation operation. Finally, we further process this intermediate output through a linear operator:
(12) $z[u/v]=tanh(\mathbf{W}_{u/v}^{z}\sum\limits_{r\in{\mathcal{R}}}z^{r}[u/v]),$
$\Gamma[u/v]=tanh(\mathbf{W}_{u/v}^{\Gamma}\sum\limits_{r\in{\mathcal{R}}}\Gamma^{r}[u/v]),$
where $\mathbf{W}_{u/v}^{z}$ and $\mathbf{W}_{u/v}^{\Gamma}$ denote the weight matrices in the nonlinear activation function $tanh(\cdot)$, $z[u/v]$ and $\Gamma[u/v]$ represent local and global embeddings of users/items after cross-view aggregation. The final representation of users and items, $e[u/v]$, is the combination of local embedding and global embedding:
(13) $e[u/v]=z[u/v]+\Gamma[u/v].$
We employ contrastive learning to enhance the complementarity between local collaborative embeddings from user-item interaction graphs and global embeddings from hypergraphs linking users and items. In this framework, different views of the same user or item form positive pairs, while divergent views form negative pairs, providing auxiliary supervision from both local and global spaces. This is formalized with a global-local contrastive loss for user representations, based on the InfoNCE framework, as follows:
(14) $\mathcal{L}_{P}^{\mathbf{U}}=\sum_{u\in\mathbf{U}}-\log\frac{exp((z[u]\cdot\Gamma[u]/\gamma))}{\sum_{u^{\prime}\in\mathbf{U}}exp((z[u]\cdot\Gamma[u^{\prime}]/\gamma))},$
$\mathcal{L}_{P}^{\mathbf{V}}=\sum_{v\in\mathbf{V}}-\log\frac{exp((z[v]\cdot\Gamma[v]/\gamma))}{\sum_{v^{\prime}\in\mathbf{V}}exp((z[v]\cdot\Gamma[v^{\prime}]/\gamma))}.$
The final global-local contrastive loss is the sum of user objective and item objective:
(15) $\mathcal{L}_{P}=\mathcal{L}_{P}^{\mathbf{U}}+\mathcal{L}_{P}^{\mathbf{V}}.$
4.5. Bilinear Decoder and Model Optimization
In line with prior work (Berg et al., 2017; Shen et al., 2021), we adopt a bilinear decoder for the purpose of reconstructing edges within the user-item graph, while considering different ratings as distinct classes. We denote the reconstructed rating matrix as $\check{\mathbf{R}}$, which effectively captures the relationships between users and items. The decoder functions by bi-linearizing the potential rating levels, generating a probability distribution, and subsequently applying a softmax function:
(16) $p(\check{\mathbf{R}}_{ij}=r)=\frac{e[i]^{\mathsf{T}}\mathbf{Q}_{r}e[j]}{\sum_{r^{\prime}=1}^{\mathcal{R}}e[i]^{\mathsf{T}}\mathbf{Q}_{r^{\prime}}e[j]},$
where $\mathbf{Q}_{r}$ is the learnable parameter matrix of rating $r$. The predicted rating matrix is calculated as:
(17) $\check{\mathbf{R}}=\sum_{r\in\mathcal{R}}r\cdot p\left(\check{\mathbf{R}}_{ij}=r\right).$
We use a bilinear decoder to frame rating prediction as a classification task, adopting Cross-Entropy (CE) loss over Mean Squared Error (MSE). While MSE is common, it often converges slowly and risks suboptimal solutions due to gradual weight updates and non-convex optimization challenges (Qi et al., 2020; Jin and Montúfar, 2023).
However, traditional CE loss overlooks the sequential and comparable nature of ratings. For example, if the true rating is 3, CE treats predictions like (0.1, 0.15, 0.3, 0.25, 0.2) and (0.05, 0.2, 0.3, 0.35, 0.1) equally, even though the latter is more accurate. To address this, we enhance CE by framing the task as multi-class classification, assigning 1 to the correct rating and a smaller value, $l_{close}$, to adjacent ratings. This adjustment ensures that predictions closer to the true rating receive a lower loss, better capturing both magnitude and order information. By adopting balanced Cross-Entropy (bCE) loss, we effectively integrate the rating’s nuances into our model, improving accuracy while maintaining a balance with $l_{close}$ typically set to 0.1.
(18) $\mathcal{L}_{main}=\frac{1}{|(i,j)|\Omega_{i,j}=1|}\sum\limits_{(i,j)|\Omega_{i,j}=1}\mathbf{bCE}(r[i,j],\hat{r}[i,j]),$
We denote the true rating label as $r[i,j]$ and the predicted rating as $\hat{r}[i,j]$ for the pair $(i,j)$. We also introduce hyperparameters $\alpha$ and $\beta$ to optimize the contrastive learning tasks. This results in the final loss function, defined as follows:
(19) $\mathcal{L}=\mathcal{L}_{main}+\alpha\mathcal{L}_{S}+\beta\mathcal{L}_{P}+\lambda\mathcal{L}_{NRR}.$
Here, we introduce $\mathcal{L}_{NRR}$ as a regularization method inspired by prior models (Berg et al., 2017; Shen et al., 2021). This regularization method is designed to account for both magnitude and order information, which aims to foster similarity in the representation of each node in the rating subgraph that is adjacent to each other.
5. Experiment
In this section, we evaluate the performance of our proposed method through extensive experiments and answer the following questions:
  • (RQ1) How does MHCL effectively perform in the matrix completion task?
  • (RQ2) How does MHCL perform compared to review and rating-based recommendation methods after matrix completion?
  • (RQ3) What are the effects of different modules in MHCL on performance?
  • (RQ4) How does MHCL mitigate long-tail distribution and data sparsity problems?
  • (RQ5) How do different hyperparameter settings affect the performance?
  • (RQ6) How does MHCL perform in terms of efficiency compared to baselines? Experiment results are in Sec. A.6.1 in Supplement.
5.1. Experimental Settings
5.1.1. Datasets
To assess the efficacy of MHCL, we conduct matrix completion and recommendation experiments on 6 publicly available datasets, including Yelp (Seo et al., 2017), Amazon (Ni et al., 2019), MovieLens-100K (ML-100K for short) (Miller et al., 2003), MovieLens-1M (ML-1M for short) (Miller et al., 2003), YahooMusic (Dror et al., 2012) and Douban (Miller et al., 2003). The summary of dataset statistics is presented in Table 1. A detailed description of datasets can be found in Sec. A.2 of the Supplement.
Table 1. Statistical summaries of datasets.
Dataset #Users #Items #Ratings density(%) Ratings%
1/(1-20) 2/(21-40) 3/(41-60) 4/(61-80) 5/(81-100)
Yelp 8423 3742 88647 0.281 5.88 9.41 15.29 37.13 32.09
Amazon 5541 3568 64706 0.330 4.31 4.65 10.49 25.55 54.98
ML-100K 943 1682 100000 6.30 6.11 11.37 27.14 34.17 21.2
ML-1M 6040 3706 1000209 4.47 5.61 10.75 26.12 34.89 22.63
YahooMusic 3000 3000 5335 0.06 1.04 11.84 19.15 19.03 48.94
Douban 3000 3000 136891 1.52 5.87 11.25 25.71 30.23 26.94
5.1.2. Baselines
To demonstrate the effectiveness, we compare our proposed MHCL with the following four categories of methods: (1) traditional methods: GRALS (Rao et al., 2015), NNMF (Dziugaite and Roy, 2015); (2) GAE-based methods: GC-MC (Berg et al., 2017), IMC-GAE (Shen et al., 2021); (3) inductive methods: PinSage (Ying et al., 2018), IGMC (Zhang and Chen, 2019), GIMC (Zhang et al., 2022), IDCF-NN (Wu et al., 2021b) and IDCF-GC (Wu et al., 2021b); (4) review-based methods: SSG (Gao et al., 2020a), RGCL (Shuai et al., 2022) and MAGCL (Wang et al., 2023). Besides, we also introduce: (5) review and rating-based recommendation methods: SGL (Wu et al., 2021a), DGCLR (Ren et al., 2022), VGCL (Yang et al., 2023), and ReHCL (Wang et al., 2024) for comparison to evaluate the recommendation performance of MHCL. A detailed description of baselines can be found in Sec. A.3 of the Supplement.
5.1.3. Evaluation Metrics and Experimental Settings
We use MSE and MAE metrics for evaluating the matrix completion task, and MRR@10 (M@10 for short) and NDCG@10 (N@10 for short) for the recommendation task. The experimental settings are detailed in the Sec. A.4 of the Supplement.
5.2. Matrix Completion (RQ1)
As shown in Table 2, MHCL consistently outperforms all baselines across datasets, with the most +9.04% notable MSE improvement on ML-100K. MHCL excels without needing extra side information or review data, making it simpler than models requiring complex preprocessing. It also surpasses review-based models on Amazon and Yelp datasets, and is more efficient than subgraph-based models, avoiding the significant time and memory costs associated with creating individual subgraphs for each rating edge, especially in large datasets. The performance highlights MHCL’s superiority in mitigating data sparsity and long-tail distribution issues.
Table 2. Matrix completion performance comparison of the proposed model and different baselines. The best results are in bold and the second-best are underlined. Marker * indicates the results are statistically significant (t-test with p-value ¡ 0.01).
Method Amazon Yelp ML-100K ML-1M YahooMusic Douban
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
GRALS 0.8297 0.6879 1.1553 0.8462 0.8930 0.7481 0.6905 0.6511 1444 51.26 0.8625 0.7304
NNMF 0.8316 0.6960 1.1607 0.8530 0.8154 0.7116 0.7106 0.6663 466.56 18.43 0.8658 0.7355
GC-MC 0.8090 0.6788 1.1508 0.8429 0.8190 0.7164 0.6922 0.6520 420.25 17.01 0.8596 0.7159
IMC-GAE 0.7921 0.6635 1.1418 0.8389 0.8046 0.7068 0.6872 0.6486 349.69 14.77 0.8533 0.7092
PinSage 0.8447 0.7032 1.1715 0.8615 0.9044 0.7522 0.8208 0.6773 524.41 20.33 0.8636 0.7641
IGMC 0.7903 0.6616 1.1481 0.8452 0.8190 0.7132 0.7344 0.6773 364.81 15.41 0.8491 0.7235
GIMC 0.7889 0.6597 1.1435 0.8422 0.8010 0.7034 0.7276 0.6724 338.56 14.52 0.8401 0.7092
IDCF-NN 0.7823 0.6536 1.1415 0.8403 0.8085 0.7059 0.6901 0.6532 342.67 15.02 0.8419 0.7110
IDCF-GC 0.7791 0.6527 1.1402 0.8391 0.8030 0.7035 0.6889 0.6501 340.01 14.75 0.8402 0.7098
[…]
MHCL 0.7610* 0.6350* 1.1289* 0.8255* 0.7790* 0.6907* 0.6790* 0.6361* 334.89* 13.89* 0.8298* 0.7019*
Traditional matrix completion methods such as GRALS and NNMF perform well on small datasets but struggle with scalability, while GNN-based models excel in this area. Although GC-MC and IMC-GAE provide fast runtimes, they converge more slowly and deliver slightly inferior performance compared to MHCL. Inductive models like IGMC and GIMC achieve competitive results but require substantial time investment, with their performance on certain datasets not matching that of MHCL. Models such as PinSage, IDCF-NN, and IDCF-GC effectively capture local patterns but fail to model higher-order dependencies and cross-rating features, which diminishes their accuracy in ordinal matrix completion tasks. Review-based models like SSG and RGCL leverage user comments to enhance embeddings but cannot function without review data. In contrast, MHCL dynamically learns hypergraph structures for each rating subgraph, capturing both local and global dependencies, and integrates these representations through cross-rating fusion. Notably, MHCL does not rely on review data yet consistently outperforms review-based methods on datasets like Amazon and Yelp, showcasing its versatility. Moreover, MHCL offers significant advantages in time complexity, making it a scalable and efficient solution for large-scale matrix completion tasks.
Table 3. Recommendation performance comparison of the proposed model and different baselines. The best results are in bold and the second-best are underlined. Marker * indicates the results are statistically significant (t-test with p-value ¡ 0.01).
Method Amazon Yelp ML-100K ML-1M YahooMusic Douban
M@10 N@10 M@10 N@10 M@10 N@10 M@10 N@10 M@10 N@10 M@10 N@10
GRALS 0.2529 0.3301 0.2705 0.3364 0.2920 0.3495 0.3921 0.4133 0.2920 0.3524 0.3147 0.3655
NNMF 0.2497 0.3299 0.2701 0.3360 0.2911 0.3487 0.3907 0.4415 0.2909 0.3510 0.3134 0.3642
GC-MC 0.2555 0.3310 0.2713 0.3377 0.2928 0.3501 0.3948 0.4453 0.2942 0.3542 0.3166 0.3670
IMC-GAE 0.2557 0.3314 0.2721 0.3387 0.2939 0.3511 0.3950 0.4454 0.2945 0.3547 0.3168 0.3671
PinSage 0.2565 0.3330 0.2740 0.3411 0.2961 0.3529 0.3958 0.4466 0.2955 0.3555 0.3180 0.3682
IGMC 0.2603 0.3366 0.2768 0.3430 0.2973 0.3550 0.3988 0.4493 0.2980 0.3583 0.3201 0.3705
GIMC 0.2637 0.3379 0.2787 0.3448 0.3002 0.3570 0.4009 0.4527 0.2996 0.3607 0.3208 0.3717
IDCF-NN 0.2791 0.3409 0.2802 0.3465 0.3011 0.3585 0.4028 0.4569 0.3018 0.3625 0.3220 0.3730
IDCF-GC 0.2825 0.3417 0.2810 0.3470 0.3013 0.3597 0.4035 0.4588 0.3027 0.3645 0.3229 0.3735
[…]
MHCL 0.3153* 0.3768* 0.3130* 0.3815* 0.3317* 0.3968* 0.4214* 0.4972* 0.3233* 0.3926* 0.3471* 0.4035*
5.3. Recommendation (RQ2)
Next, we evaluate the recommendation performance by comparing MHCL with the latest recommendation baselines which consider user rating matrices. Expect baselines in Table 2, we add the review and rating-based recommendation baselines (SGL, DGCLR, VGCL, ReHCL). Results are reported in Table 3, MHCL consistently outperforms all advanced baselines across datasets, with the most +10.31% notable N@10 improvement on ML-100K and average +6.65% improvement on M@10 and N@10 across all datasets. The performance highlights MHCL’s superiority in recommendation tasks.
In recommendation tasks, many matrix-completion methods integrate user rating matrices into node representations to effectively capture user preferences and enhance overall recommendation performance. For instance, on Amazon and Yelp, which include user reviews, the current state-of-the-art method, ReHCL, leverages both the rating matrix and the textual reviews provided by users. In contrast, our method, MHCL, achieves an average performance improvement of +4.14% over ReHCL without incorporating any review data, relying solely on the rating matrix. This outcome highlights the effectiveness and necessity of our matrix completion algorithm design, demonstrating that a well-structured rating matrix can be sufficient for improving recommendation quality. Furthermore, when analyzing four datasets that do not contain reviews, we observe a maximum performance improvement of over 10% in both M@10 and N@10 metrics compared to IDCF-GC, the leading matrix completion method among the baselines. This finding suggests that the rating matrices generated through MHCL not only capture more nuanced user preference information but also maintain their effectiveness across diverse datasets. Overall, MHCL’s ability to deliver superior performance without relying on review data underscores its potential as a robust solution for matrix completion in recommendation systems.
5.4. Ablation Study (RQ3)
To evaluate the effectiveness of each component of our model, we have further conducted an ablation study on different MHCL variations. We report the results of the ablation study on five datasets in Table 4. Specifically, we generate the following variants:
  • w/o C-R - which removes the multi-channel cross-rating contrastive learning and corresponds to the $\alpha\mathcal{L}_{S}$ part of the final loss function.
  • w/o Glo - which removes the adaptive hypergraph structure learning module and only learns the local node embeddings. This variant corresponds to the $\beta\mathcal{L}_{P}$ part in the final loss function.
  • w/o CL - which removes all contrastive learning mechanisms.
  • w/o CRF - which removes the cross-rating representation fusing module.
  • w/o bCE - which uses the traditional cross-entropy loss instead of the balanced CE loss that captures the nuances of ratings.
  • w/o NRR - which removes the $\mathcal{L}_{NRR}$ in the final loss function.
  • with OCE - which uses the ordinal cross-entropy loss (OCE) (Shi et al., 2023) instead of the bCE loss that captures the nuances of ratings.
Our results yield several key insights. First, removing the cross-ratings representation fusion (CRF) module (w/o CRF) leads to performance degradation across all datasets, underscoring its critical importance in capturing complex user interactions. Similarly, omitting the multi-channel cross-ratings module and global hypergraph learning (w/o C-R and w/o Glo) also results in reduced performance. Notably, the cross-ratings module significantly enhances results for Amazon, where the rating distribution is imbalanced, while hypergraph learning offers particular benefits for denser datasets such as ML-100K and ML-1M. The variant without contrastive learning (w/o CL) demonstrates the poorest performance, highlighting the crucial role of contrastive learning in enriching node representations and improving model robustness. Additionally, removing the balanced cross-entropy loss (w/o bCE) adversely affects performance, further confirming its effectiveness in capturing correlations between closely-rated scores. Furthermore, our analysis indicates that the cross-ratings module (C-R) has the most substantial impact on the loss function, whereas node relation regularization (NRR) exhibits the least effect, aligning with our parameter sensitivity findings. Finally, while the ordinal cross-entropy loss (with OCE) performs similarly to the balanced cross-entropy loss, the latter is preferred due to its slight performance edge and superior computational efficiency, making it a more practical choice for implementation.
Table 4. Performance comparison of different variants (MSE).
Dataset Amazon YELP ML-100K ML-1M YahooMusic
MHCL (w/o C-R) 0.7839 1.1320 0.7856 0.6842 336.72
MHCL (w/o Glo) 0.7703 1.1354 0.7899 0.6874 340.53
MHCL (w/o CL) 0.7884 1.1392 0.7910 0.6905 350.77
MHCL (w/o CRF) 0.7701 1.1320 0.7857 0.6857 342.13
MHCL (w/o bCE) 0.7722 1.1343 0.7836 0.6851 339.85
MHCL (w/o NRR) 0.7689 1.1328 0.7845 0.6869 340.05
MHCL (with OCE) 0.7651 1.1298 0.7796 0.6799 335.95
MHCL 0.7610 1.1289 0.7790 0.6790 334.89
5.5. Impact of imbalanced data distribution (RQ4)
5.5.1. Impact of Imbalanced User Interactions
We evaluate MHCL’s performance across users with varying interaction sparsity by dividing them into three groups: Inactive (bottom 80% by interaction frequency), Normal (middle 15%), and Active (top 5%). We then compare MHCL with two leading baselines on two datasets, with results shown in Figures LABEL:fig2.1 and LABEL:fig2.2. MHCL consistently outperforms the baselines, with the largest gains observed among Inactive users. This improvement is driven by InfoNCE loss, which mitigates popularity bias by enhancing representation uniformity.
5.5.2. Impact of Imbalanced Rating Interactions
Experimental results according to different rating categories are shown in Figures LABEL:fig2.3 and LABEL:fig2.4. The findings highlight significant improvements in long-tail ratings (‘1’ and ‘2’) without sacrificing performance in the head classes. This underscores the effectiveness of InfoNCE in reshaping embedding distributions to better capture long-tail characteristics. The impact is especially notable in imbalanced datasets like Amazon, where ratings ‘1’ and ‘2’ constitute less than 10%, making MHCL’s enhancements particularly striking.
5.6. Parameter Sensitivity (RQ5)
5.6.1. Effect of Hyperparameter $\alpha$, $\beta$ and $\lambda$.
We analyze the sensitivity of parameters $\alpha$, $\beta$, and $\lambda$ on contrastive learning (CL) losses using the Amazon dataset (Fig. LABEL:alpha_amazon, LABEL:beta_amazon, LABEL:lambda_amazon). Increasing these parameters from 0 to 0.01 enhances MHCL’s performance, but larger values can impair the main loss function. The optimal range of 0.01 to 0.05 highlights the need to balance CL with the main loss for optimal results.
5.6.2. Effect of Hyperedge Number
We explore the optimal number of hyperedges, varying them from 8 to 128 on the Amazon dataset and from 64 to 320 on ML-100K (Fig. LABEL:r_ml). Fig. LABEL:r_amazon shows that Amazon achieves peak performance with 16 hyperedges, while higher numbers lead to noise and reduced performance. Sparser datasets like Amazon benefit from fewer hyperedges to avoid overfitting.
6. Conclusion
In this paper, we introduce a novel Multi-Channel Hypergraph Contrastive Learning (MHCL) framework for matrix completion. First, we use a hypergraph learner to construct hypergraph structures corresponding to different rating subgraphs, capturing high-order dependencies and obtaining global node embeddings through HGNN. Second, we enhance the representation of ratings by encouraging alignment between adjacent ratings. Additionally, we employ global-local contrastive learning to simultaneously characterize local and global embeddings in a joint space. Experimental results on five benchmark datasets demonstrate the superiority and effectiveness of our MHCL.
Appendix A Supplement
A.1. Notations
Key notations used in the paper are summarized in Table 5.
Table 5. The summary of the main used notations.
Notation Definition
$\mathbf{U},\mathbf{V}$ Two types of node sets (users and items)
$M$, $N$ The number of users and items
$\mathbf{R}$ The entire user-item rating matrix
$\mathcal{R}$ The set of all ratings
$G$ Multiplex user-item bipartite graph
$\mathbf{E}$ The entire learnable parameter matrix
$K$ The number of hyperedges
$x^{r}$ Local embeddings of nodes within the $r$-th rating subgraph
$\mathcal{N}^{r}(i)$ The neighbors of node $i$ in $r$-th rating subgraph
$\mathbf{H}_{u}^{r},\mathbf{H}_{i}^{r}$ Adaptive hypergraph of nodes on the $r$-th rating subgraph
$h_{u}^{r},h_{i}^{r}$ Global embeddings of nodes in the $r$-th rating subgraph
$z,\Gamma$ Local and global embeddings after cross-view aggregation
$e$ The final representation of nodes
$d$ Dimension of the embedding
$\mathbf{W},\mathbf{q}$ Learnable weight matrix
$\mathcal{N}_{r}$ The two ratings that are adjacent to the rating $r$
A.2. Detailed Dataset Description
For YahooMusic dataset, we utilize preprocessed subsets of this dataset as provided by (Monti et al., 2017). Amazon and Yelp datasets are distinct in that they include not only user-item ratings but also detailed review information. These datasets encompass four primary data components: users, items, ratings, and review documents. We focus on the ‘Digital Music’ domain within the Amazon dataset, specifically the Amazon-5core dataset in our experiments, where ‘5-core’ indicates that each user or item has a minimum of five reviews. Yelp dataset originates from the Yelp Business Rating Prediction Challenge 2013 dataset, which comprises restaurant reviews from the Phoenix, AZ metropolitan area. Following the preprocessing steps outlined in (Shuai et al., 2022), we establish a 5-core dataset from the raw Yelp data.
A.3. Baselines
We compare our MHCL with the following 12 baselines, which can be deeply divided into four categories: Traditional Method:
  • GRALS (Rao et al., 2015) is a graph-regularized matrix completion algorithm, which has formulated and derived an exceptionally efficient alternating minimization scheme based on conjugate gradients. The scheme excels in solving optimizations involving over 55 million observations.
  • NNMF (Dziugaite and Roy, 2015) is a classical matrix factorization model which uses the user-item ratings (or interactions) only as the target value of their objective function. It explores the possibility of substituting the inner product with a learned function, concurrently with the learning process of latent feature vectors.
GAE-based Methods:
  • GC-MC (Berg et al., 2017) regards rating prediction as link prediction on the user-item bipartite graph and adopts Graph Autoencoder (GAE) to encode user and item embeddings, and proposes a graph auto-encoder framework which can utilize differentiable message passing on the bipartite interaction graph.
  • IMC-GAE (Shen et al., 2021) is an inductive GAE-based matrix completion method that leverages GAE to acquire personalized user or item representations for tailored recommendations while also capturing local graph patterns for inductive matrix completion.
Inductive Methods:
  • PinSage (Ying et al., 2018) combines highly effective random walk and graph convolution to represent node embeddings that incorporate both node features and graph structure. It also introduces an innovative training strategy, which progressively incorporates more challenging training examples to enhance the model’s robustness and convergence.
  • IGMC (Zhang and Chen, 2019) employs a GNN-based model that is exclusively trained on 1-hop subgraphs centered around (user, item) pairs derived from the rating matrix. It exhibits inductive capabilities, learning local graph patterns to generalize to new ones for inductive matrix completion.
  • GIMC (Zhang et al., 2022) stands as the pioneer in defining continuous explicit feedback prediction within a non-Euclidean space by introducing hyperbolic geometry and a unified message-passing scheme to reflect node interaction and represent edge embeddings for improving inductive matrix completion for rating prediction.
  • IDCF-NN (Wu et al., 2021b) is an inductive collaborative filtering method that adopts the traditional matrix factorization method to factorize the rating matrix of a set of key users to obtain meta-latent factors.
  • IDCF-GC (Wu et al., 2021b) is an inductive collaborative filtering method, which is an attention-based structured learning method that estimates the hidden relationship from query to key users and learns to use meta-latents to inductively compute the embedding of query users through neural information transfer.
Review-based Methods:
  • SSG (Gao et al., 2020a) introduces a multi-view approach to enhance existing single-view methods, incorporating two additional perspectives: sequence and graph, for leveraging reviews. Specifically, three-way encoder architecture captures long-term (set), short-term (sequence), and collaborative (graph) user-item features for recommendations.
  • RGCL (Shuai et al., 2022) is a novel graph-based contrastive learning framework that firstly creates a review-aware user-item graph with edge features enriched by review content, and then introduces two additional contrastive learning tasks to provide self-supervised signals.
  • MAGCL (Wang et al., 2023) is a review-based recommendation method that captures fine-grained semantic information in user reviews while considering the intrinsic correlation between ratings and reviews, and learns higher-order structure-aware and self-discriminative node embeddings through multi-view representation learning, graph contrastive learning, and multi-task learning mechanisms.
Review and Rating-based Recommendations
  • SGL (Wu et al., 2021a) enhances GCN-based recommendation by adding a self-supervised task. It generates multiple node views using node dropout, edge dropout, and random walk, maximizing agreement between views of the same node. Implemented on LightGCN, SGL improves recommendation accuracy, especially for long-tail items, and increases robustness to noise.
  • DGCLR (Ren et al., 2022) models user-item interactions through latent factors derived from reviews and user-item graphs. A factorized message-passing mechanism learns disentangled representations, while an attention mechanism combines predictions from multiple factors. Factor-wise contrastive learning mitigates sparsity and improves factor-specific interaction modeling.
  • VGCL (Yang et al., 2023) estimates a Gaussian distribution for each node, generating contrastive views through sampling. These views reconstruct the input graph without distortion, and node-specific variance adjusts contrastive loss. A twofold contrastive learning approach ensures consistency at both the node and cluster levels.
  • ReHCL (Wang et al., 2024) constructs topic and semantic graphs to mine review relations and employs cross-view contrastive learning to enhance node representations. A neighbor-based positive sampling captures similarities between views, reducing noise. Additionally, cross-modal contrastive learning aligns rating and review representations. These modes form a hierarchical contrastive learning framework to improve recommendation performance.
It is worth noting that certain methods within these groups require specific preprocessing steps before model training. For instance, RGCL employs BERT-Whitening (Su et al., 2021) to convert textual comments into vector representations, while both IGMC and GIMC mandate the extraction of subgraphs for all edges within the dataset to operate effectively. For most of the baselines, we split our training, validation, and test sets in the ratio of 8:1:1 following the same settings in recent studies (Zhang et al., 2022; Shen et al., 2021).
A.4. Parameter Setting
We employ the Adam optimizer for the entirety of model training, initializing all trainable parameters using the Xavier method. We systematically explore several key hyperparameters: The number of layers is varied within the range $\{1,2,...,5\}$. The hyperparameters $\alpha$, $\beta$ and $\lambda$ in a range of $\{0.001,0.005,0.01,0.05,0.1\}$. In the embedding layer, we consider different vector sizes, selecting from {30, 60, …, 1200}. The number of hyperedges is searched in $\{8,16,32,128,256,512\}$. Additionally, we implement early stopping to mitigate the risk of overfitting. The decay ratio, denoted as $\theta$, is set to 0.5 to further fine-tune training dynamics. The complete model implementation is carried out using the Deep Graph Library (Wang, 2019) and PyTorch, harnessing the computational power of an Intel Xeon Gold 5320T CPU (2.3GHz) with two NVIDIA RTX 3090 GPUs.
Specifically, in the Recommendation section, we utilize the following loss function to evaluate the performance of our MHCL:
(20) $\mathcal{L}=\sum\limits_{(u,{v_{+}},{v_{-}})\in\mathcal{O}}{-\ln(Sigmoid(\hat{y}_{u,{v_{+}}}-\hat{y}_{u,{v_{-}}}))}+\lambda||\Theta|{|^{2}}.$
where $\hat{y}_{u,{v_{+}}}=e_{u}[i]^{\mathsf{T}}e_{v_{+}}[i]$, $(u,{v_{+}})\in\mathcal{O}_{+},(u,{v_{-}})\in\mathcal{O}_{-}$, $\mathcal{O}_{+}$ and $\mathcal{O}_{-}$ respectively denote the observed existent interaction and unobserved interaction, and $\mathcal{O}=\mathcal{O}_{+}+\mathcal{O}_{-}$.
A.5. Time Complexity Analysis
The proposed MHCL consists of four key components: subgraph partition and graph convolution module, multi-channel cross-rating contrastive learning module, adaptive hypergraph structure learning module, and attention-based cross-view aggregation module. For the subgraph partition and graph convolution module, the time complexity is $O((M+N+3)\times\mathcal{|R|}\times L\times d)$, where $\mathcal{|R|}$ represents the number of rating categories in the user-item entire matrix, $M$ and $N$ represent the number of users and items, $L$ represents the number of GCN message propagation layer. The time complexity of multi-channel cross-rating contrastive learning module is $O((M+N)\times d^{2}\times\mathcal{|R|}+(M+N)\times d\times\mathcal{|R|})$. The time complexity of the adaptive hypergraph structure learning module is $O((M+N)^{2}\times d)$. For the attention-based cross-view aggregation module, the time complexity is $O(4(\mathcal{|R|}-1)^{2}\times d)$. Conclusively, the total time complexity of our MHCL is $O((M+N)^{2}\times d+(M+N)\times\mathcal{|R|}\times d\times(d+L+1)+4(\mathcal{|R|}-1)^{2}\times d+3\mathcal{|R|}\times L\times d)$.
A.6. Additional Experiments
A.6.1. Efficiency Evaluation (RQ6)
Table 6. Running time for each epoch and total time cost in different datasets.
Method ML-100K Yelp Amazon YahooMusic
Epoch Total Epoch Total Epoch Total Epoch Total
GC-MC 0.15 2.25 1.07 16.05 0.45 6.75 0.35 5.25
IMC-GAE 0.23 3.45 1.17 17.55 0.65 9.75 0.37 5.55
IGMC 706 7060 362 3620 267 2670 19.3 193
GIMC 600 6000 231 2310 143 1430 12.7 127
IDCF-NN 0.56 5.6 2.67 26.7 1.45 14.5 0.97 9.7
IDCF-GC 0.60 6.0 2.85 28.5 1.59 15.9 1.01 10.1
SSG 0.76 11.4 3.01 45.15 1.75 26.25 0.93 13.95
RGCL 0.46 6.9 2.41 36.15 1.14 17.1 0.79 11.85
MAGCL 0.75 11.25 2.97 44.55 1.79 26.85 1.01 15.15
MHCL 0.45 4.5 2.42 24.2 1.13 11.3 0.75 7.5
We proceed to assess the actual running time, denoted as $t$, expended by the compared methods for a single epoch, and model total time cost which counts the total time for our method to reach convergence on the four datasets. These results are depicted in Table 6, we scrutinize the inference time of MHCL in comparison to GC-MC, IMC-GAE, IGMC, GIMC, IDCF-NN, IDCF-GC, SSG, RGCL, and MAGCL across four datasets. Notably, the two inductive models, IGMC and GIMC, incur a higher time overhead. This is primarily due to the necessity for these models to extract a local subgraph for all rating edges and subsequently relabel the nodes within each subgraph. As improved inductive approaches, IDCF-NN and IDCF-GC provide a huge improvement in time cost compared to IGMC. According to (Shen et al., 2021), for a graph with $N$ nodes and $M$ ratings, GAE-based models such as GC-MC and IMC-GAE require applying the GNN $O(N)$ times to compute an embedding for each node. In contrast, IGMC and GIMC demand applying the GNN $O(M)$ times for all ratings. When the ratio of $M$ to $N$ is significantly large ($M\gg N$), IGMC exhibits worse time complexity compared to GAE-based models. Meanwhile, GIMC, by introducing edge embeddings to capture the semantic properties of different ratings rather than treating edges as propagation channels within a conventional message-passing scheme, boasts a lower time complexity than IGMC. For review-based methods such as SSG, RGCL and MAGCL, their time complexity is significantly higher than that of our proposed MHCL due to the additional comment information they need to process, and the actual results on several datasets are slightly inferior to those of MHCL.
Note that although our MHCL consumes more total time than GC-MC and IMC-GAE, in practice our method converges faster than the above two methods, which keeps the actual total consumed time all in the same order of magnitude, and our method does not lag behind much, and MHCL achieves a better performance.
Furthermore, owing to the construction of hypergraphs and the influence of contrastive learning, MHCL demonstrates slightly higher time complexity than IMC-GAE. In particular, the incorporation of hypergraph structure learning and contrastive learning contributes to the additional computational overhead in MHCL.
A.6.2. Effect of Hyperparameter $\alpha$, $\beta$ and $\lambda$.
We evaluate the impact of the contrastive learning losses by conducting parameter sensitivity experiments with $\alpha$, $\beta$ and $\lambda$ on ML-100K dataset, using values from {0.001, 0.005, 0.01, 0.05, 0.1}. Results for $\alpha$, $\beta$ and $\lambda$ are shown in Figures LABEL:alpha_ml,  LABEL:beta_ml and  LABEL:lambda_ml. We find that increasing $\alpha$ and $\beta$ from 0 to 0.01 improves MHCL’s performance on ML-100K dataset, indicating that contrastive learning enhances node representation learning. However, excessively large values diminish the model’s main loss function, negatively impacting performance. The optimal range for $\alpha$ and $\beta$ is between 0.01 and 0.05, highlighting the need to balance the main loss with contrastive learning for optimal results. Moreover, despite having fewer users and items, MHCL requires more hyperedges to capture complex high-order correlations on ML-100K dataset (as shown in Fig. LABEL:r_ml).
A.6.3. Effect of Convolution Layers.
To assess MHCL’s performance with varying depths, we experiment with layer counts from 0 to 5 on the Amazon and ML-100K datasets. As shown in Table 7, performance improves with increasing depth, peaking at 3-4 layers, after which it declines. This suggests that MHCL effectively captures high-order collaborative signals while mitigating over-smoothing. Notably, MHCL accommodates deeper layers compared to other GNN models like RGCL (Shuai et al., 2022) (1 layer) and IMC-GAE (Shen et al., 2021) (2 layers), underscoring its ability to model complex relationships in data.
Table 7. MSE comparison with different layers
Number of Layers MSE on Amazon MSE on ML-100K
1 Layer 0.7825 0.7940
2 Layers 0.7691 0.7852
3 Layers 0.7610 0.7790
4 Layers 0.7657 0.7833
5 Layers 0.7702 0.7871